Consider a binomial distribution random variable X with 20 trials (n=20) and p=0.45 (the probability of sucess). What is the average?

Answer by stanbon(75887)

Consider a binomial distribution random variable X with 20 trials (n=20) and p=0.45 (the probability of sucess). What is the average?
-------
average = np = 20*0.45 = 9
==============================
Cheers,
Stan H.
===========